If boron hydride, B4H10, is treated with pure oxygen, it burns to give B2O3 and H2O

2 B4H10 (g) + 11 O2 (g) → 4 B2O3 (s) + 10 H2O (g)

If a 0.05 g sample of the boron hydride burns completely in oxygen, what will be the pressure of the gaseous water in a 4.25 L flask at 30.0oC?


1
Expert's answer
2021-04-19T05:10:58-0400

Write the balanced equation and you will see that 1 mole of B4H10 will yield 5 moles of H2O


Calculate the moles of B4H10 present in 0.050 g. (0.050 g / 53.3 g/mol = 9.378 × 10-4 mol)

Multiply the mole of B4H10 by 5 to get n (moles of H2O = 5 × 9.378x10-4 = 0.00469 mol)

PV=nRT

P (atm) ={0.00469 × 0.082057 × (273+30) }/ 4.25 L

P = 0.027 atm (21 torr)